OpenStudy (lalala):

Which equation is the perpendicular bisector of the line segment with endpoints (-2, 4) and (6, 8)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

first you need the slope of this line which is \[m=\frac{8-4}{6-(-2)}=\frac{4}{8}=\frac{1}{2}\]

OpenStudy (anonymous):

since \[m=\frac{1}{2}\] this means the perpendicular line has slope \[m=-2\] the negative reciprocal now we need to find a point on that line

OpenStudy (anonymous):

the midpoint between (-2,4) and (6,8) is \[(\frac{-2+6}{2},\frac{4+8}{2})=(2,6)\]

OpenStudy (anonymous):

use point slope formula to get \[y-6=-2(x-2)\] \[y-6=-2x+4\] \[y=-2x+10\]